OpenStudy (anonymous):

Part 1: Solve |2x-9|+4 less than or equal to 7. You must show all work. Use >= and <= to represent the inequalities. Part 2: Use complete sentences to describe the graph in words

OpenStudy (mertsj):

Subtract 4 from both sides.

OpenStudy (ash2326):

|2x-9| +4 <=7 let's first assume (2x-9) is greater than equal to zero, when we get the solution it'll verify if we are wrong or right is 2x-9>=0 x>=9/2 x>=4.5 now we have 2x-9+4<=7 2x<=12 x<=6 this lies in the region of x>=4.5 taking intersection of the two ranges we get 4.5<=x<=6 now let's assume 2x-9 <0 we get x<4.5 now modulus will open by a negative sign |2x-9|+4<=7 -2x+9+4<=7 13-7<=2x or 2x>=6 x>=3 and we have a pre condition x<4.5 so intersection of these two gives 3<=x<4.5 so complete solution is x belongs to [3,4.5) U [4.5,6] which gives x is [3,6]

OpenStudy (anonymous):

Both sides of what? I know you took all your time to type that, but I dont understand it.

OpenStudy (anonymous):